• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 00:02:32 on May 5,2000

    Entry number 40511

    keyword=Shift Summary

    We started the shift by finishing the arc energy measurement. At 5:00 pm we
    started taken ep data with the beam in dispersive mode, we then switched to
    acromatic mode and took a normal ep measurement. Starting at 7:30 we began
    taking 4He data again.

    At 10:30 we tried to take a BPM pedestal run, but the data didn't look good.
    We believe the problem is with how MCC is setting the forced gain to zero,
    though we are not sure.

    Files 2033 and 2034 were for the BPM pedestal measurement.

    Files Events
    2031 250k
    2032 250k
    2035 160k Hadron Q1 Error at end of run, data should be OK.
    -----------------------
    Total 660 k